Electric potential at the surface of an atomic nucleus z=50 of radius 9×10^-15
 

Grade:12

3 Answers

Arun
4 years ago
Note that in a nucleus there are only one entity that contains charge i.e proton. Neutron also resides in nucleus but it is neutral. 

since Z = atomic number = number of protons each having electronic charge e = 1.6*10^-19C  

V = (k)*(Q)/R ------------------[ potential at the surface of nucleus] 

V = (k)*(Z*e)/R ------------------[ Total charge = (number of protons)*e ] 

k = 9*10^9Nm^2/C^2 , Z = 50 , e = 1.6*10^-19C , r = 9*10^-15m 

V = (9*10^9*(50)*(1.6)*10^-19)/(9*10^-15) 

V = (9*50*1.6*10^9-19)/(9*10^-15) 

V = 9*80*10^-10/9*10^-15 

V = 80*(10^-10)/(10^-15) 

V = 80*[10^(-10-(-15))] 

V = 80*10^5 

so V = 8*10^6 volts
Khimraj
4 years ago
V= KQ/r = KZe/r
K = 9*109
Z=50
e = 1.6*10-19
r = 9*10-15
so V = 8*106 volts
….............................................................
praga theesh
4 years ago
V = (k)*(Q)/R ------------------[ potential at the surface of nucleus] 

V = (k)*(Z*e)/R ------------------[ Total charge = (number of protons)*e ] 

k = 9*10^9Nm^2/C^2 , Z = 50 , e = 1.6*10^-19C , r = 9*10^-15m 

V = (9*10^9*(50)*(1.6)*10^-19)/(9*10^-15) 

V = (9*50*1.6*10^9-19)/(9*10^-15) 

V = 9*80*10^-10/9*10^-15 

V = 80*(10^-10)/(10^-15) 

V = 80*[10^(-10-(-15))] 

V = 80*10^5 
so v =80*10^6